OpenStudy (anonymous):

9a^2/3a+5b-25b^2/3a+5b simplify to lowest terms

OpenStudy (anonymous):

need some parentheses here, or a screenshot

OpenStudy (anonymous):

9a^2/3a+5b - 25b^2/3a+5b

OpenStudy (anonymous):

\[\frac{9a^2}{3a+5b}-\frac{25b^2}{3a+5b}=\frac{9a^2-25b^2}{3a+5b}\] is a start then factor

OpenStudy (anonymous):

numerator is the difference of two squares, you get \[\frac{(3a+5b)(3a-5b)}{3a+5b}\] i will let you cancel on your own

OpenStudy (anonymous):

doesnt that all cancel??

OpenStudy (anonymous):

yes, you get only \(3a-5b\) as your answer